OpenStudy (anonymous):

from a standard deck of cards, what is the probability that you choose a red card and then choosing a 3 assuming you replaced the first card

OpenStudy (anonymous):

half of the cards are red

OpenStudy (anonymous):

\(\frac{1}{13}\) of the cards are 3's multiply \(\frac{1}{2}\times \frac{1}{13}\)